OpenStudy (anonymous):

Cos 7pi/6 on unit circle

OpenStudy (lgbasallote):

are you familiar with the trigonometric formula \[cos (a + b) = \cos (a)\cos (b) - \sin (a)\sin (b)\]because you can solve this using that formula...